# HugAgent: Evaluating LLMs in Simulating Individual-Level Human Reasoning on Open-Ended Tasks

### v2

@@ Abstract
Simulating human reasoning in open-ended tasks has been a long-standing aspiration in AI and cognitive science. While large language models now approximate human responses at scale, they remain tuned to population-level consensus, often erasing the individuality of reasoning styles and belief trajectories. To advance the vision of more human-like reasoning in machines, we introduce HugAgent (Human-Grounded Agent Benchmark), a benchmark for average to individual reasoning adaptation. The task is to predict how a specific person would reason and update their beliefs in novel scenarios, given partial evidence of their past views. HugAgent adopts a dual-track design: a synthetic track for scale and systematic stress tests, and a human track for ecologically valid, “out-loud” reasoning data. This design enables scalable, reproducible evaluation of intra-agent fidelity: whether models can capture not just what people believe, but how their reasoning evolves. Experiments with state-of-the-art LLMs reveal persistent adaptation gaps, positioning HugAgent as the first extensible benchmark for aligning machine reasoning with the individuality of human thought. Our benchmark and chatbot are open-sourced as HugAgent and TraceYourThinking.
@@ 1 Introduction
Large language models (LLMs) are increasingly used as social simulators—to role-play individuals, build digital twins, and generate silicon samples for testing social and policy ideas(Park et al., 2024; Argyle et al., 2023; Xie et al., 2024a; Jiang et al., 2022). These systems promise scalability and accessibility: instead of recruiting thousands of people, researchers and practitioners can use LLMs to approximate human perspectives at scale. Yet because LLMs are pretrained on population-level corpora, they tend to collapse into an “average voice,” capturing consensus patterns while erasing the individuality of personal histories, beliefs, and reasoning styles(Wang et al., 2025; Santurkar et al., 2023; Durmus et al., 2023).
@@
Current benchmarks fail to capture this ability, across three key dimensions. 1 Intra-agent vs. inter-agent fidelity. Existing pluralistic alignment benchmarks probe group dynamics and social influence(Sorensen et al., 2024), but neglect whether models can faithfully reproduce reasoning within a single agent. 2 Reasoning traces vs. behavioral outcomes. Large-scale “digital twin” datasets such as Agent Bank (Park et al., 2024) and Twin-2K-500 (Toubia et al., 2025) primarily assess static behavioral outcomes, but not the evolving reasoning trajectories of a single individual, which are essential for credible social simulation (Li et al., 2025). 3 Open-ended vs. vignettes. Commonsense and social reasoning benchmarks (e.g., SocialIQA, ATOMIC) often reduce diverse answers to a single ground truth (Sap et al., 2019a; c). Opinion-oriented datasets likewise emphasize aggregate patterns over individual variation (Argyle et al., 2023; Santurkar et al., 2023). Theory-of-Mind style tests typically rely on short vignettes with designer labels (Wimmer & Perner, 1983; Chen et al., 2024), which limits ecological validity and overlooks first-person reasoning traces that could serve as a richer gold standard (Ying et al., 2025).
@@
Motivated by these gaps, we introduce HugAgent, a benchmark that targets intra-agent fidelity by operationalizing average-to-individual reasoning adaptation as a measurable task. For Dimension 1, HugAgent shifts the granularity from inter-agent to intra-agent fidelity: given a person’s profile and reasoning history, a model must predict both their current belief state and how it would evolve when presented with new counterfactual evidence. In Dimension 2, HugAgent advances beyond static outcomes toward reasoning trajectories. It collects first-person, out-loud self-reports as gold-standard reasoning traces. These traces offer a deeper target for prediction than the choice outcomes or survey responses typically captured in lab experiments. To address Dimension 3, instead of relying on vignette-style benchmarks, HugAgent builds evaluation around open-ended contexts. We curated real-world topics, beginning with socially and politically controversial issues that introduce inherent conflicts. Through sustained follow-up questions, the benchmark probes participants’ deliberate, System 2 style reasoning(Kahneman, 2011; Evans & Stanovich, 2013), transforming the dataset from toy settings into complex, open-ended domains. For a broader discussion of prior work on personalization, social reasoning, and user modeling in LLMs, we refer readers to Appendix A.
To build such a benchmark at scale while retaining ecological validity(i.e., the extent to which findings reflect reasoning as it occurs in real-world contexts), HugAgent combines two complementary tracks. A synthetic track provides large, controlled datasets where belief shifts and reasoning paths can be systematically manipulated(Yukhymenko et al., 2024; Xie et al., 2024b). A human-grounded track applies the same protocol to real participants, yielding data anchored in individual variation(Srivastava et al., 2023).
@@
  • How well do today’s models perform? We introduce HugAgent, a dual-track benchmark (synthetic + human) that evaluates both Belief-State Inference and Belief-Dynamics Update. Initial experiments with state-of-the-art LLMs provide baseline results and reveal adaptation gaps. https://anonymous.4open.science/r/HugAgent
@@ 2.1 Formalization
Tasks. (i) Belief-State Inference: infer stance/factor polarity from $\mathcal{C}_{i}$. (ii) Belief Dynamics Update: predict stance shifts $\widehat{\Delta\mathbf{s}}_{i}$ given $(\mathcal{C}_{i},\mathcal{I})$. Metrics include accuracy, mean absolute error (MAE), and rank correlation.
@@ 2.2 Theoretical Anchors: Probabilistic and Causal Perspectives
We use normative models as anchors rather than assumptions. (1) Bayesian / PLoT. Idealized revision follows Bayesian conditioning $b_{i}^{\prime}(\mathbf{s})\propto b_{i}(\mathbf{s})\,p(\mathcal{I}\mid\mathbf{s})$, interpreting language as probabilistic evidence over latent stances. We also draw on the Probabilistic Language of Thought (PLoT) framework(Goodman et al., 2015), which extends Bayesian inference to compositional linguistic structures. (2) Structural Causal Models (SCM). Interventions act as $do(\mathcal{I})$ on a causal graph of values/reasons, yielding counterfactual belief shifts $\mathbb{E}[\mathbf{s}\mid do(\mathcal{I})]$. We additionally represent a person’s value–reason structure as a signed directed graph $G_{i}$; we hypothesize that similarity between such graphs (e.g., via graph edit distance or learned embeddings) may predict cross-domain transfer, motivating our Hypothesis H2 and left as future work. Human reasoning deviates from these ideals; the anchors provide principled baselines for analysis.
@@ 3 HugAgent Benchmark
Figure 2: Two benchmark tasks. Task 1 (Belief-State Inference) infers stance and reasons from prior context; Task 2 (Belief-Dynamic Update) predicts stance shifts and its reasoning under new evidence.
@@ 3.1 Design Principles
  • Open-ended but deeper. Emphasize depth over breadth: semi-structured dialogue with targeted follow-ups surfaces individuality while avoiding over-scaffolding; ground truth comes from self-reports (Kvale, 1996; Srivastava et al., 2023; Park et al., 2024).
  • Two observable proxies. We evaluate (i) belief-state inference and (ii) belief-dynamics update—tractable targets that avoid requiring exact trace imitation; cf. proxy-label benchmarks (Geva et al., 2021; Ho et al., 2022; Guerdan et al., 2023).
@@
  • Human ceiling. Test–retest reliability defines the upper bound, aligning with psychology standards (Nunnally & Bernstein, 1994; Cronbach, 1970) and recent large-scale simulations (Toubia et al., 2025; Park et al., 2024).
@@ 3.2 Task Definition
We instantiate two tasks: (i) Belief-State Inference: predict $(s_{t},\mathbf{w}_{t})$ from prior responses. (ii) Belief Dynamics Update: predict $(s_{t+1},\mathbf{w}_{t+1})$ given $(b_{t},e)$. Figure 2 shows concrete examples of both tasks in HugAgent.
Figure 3: HugAgent benchmark pipeline. Inputs (demographics, questionnaires, and transcripts) flow through two components: a questionnaire that provides demographic anchors, stance baselines, and counterfactual updates, and a semi-structured chatbot that elicits individualized reasoning. Together these elements define two benchmark tasks: (i) belief-state inference—recovering stance and factor polarity from context, and (ii) belief-dynamic update—predicting stance shifts and reweighting under new evidence. (A) The chatbot maintains a causal belief network of factors, used to identify the most critical nodes and edges for follow-up. (B) A question generator derives targeted, context-specific probes from this network to structure the dialogue.
@@ 3.3 Building HugAgent: Scalable Elicitation of Individual Reasoning
HugAgent is built through a two–stage pipeline (Figure 3). The survey stage collects demographics, a baseline stance ($s_{t}$, 1–10), reason weights ($\mathbf{w}_{t}$, 1–5), and counterfactual interventions with updated stances and reasons. These structured responses provide gold labels for Belief Dynamics Update and serve as anchors for aligning free text to factors.
The chatbot stage elicits 8–20 question–answer pairs through semi-strutctured interview, combining open-ended elaborations (Context QAs) with concise polarity judgments (GT QAs). This setup captures both participants’ reasoning styles and their explicit factor-level commitments. A single transcript thereby supports both benchmark tasks: Belief-State Inference (using Context and GT QAs) and Belief Dynamics Update (using survey anchors and interventions). Survey-provided updates are never revealed in dialogue, preventing leakage.
@@ 3.4 Dataset Statistics
From over 120 participants, we retained 54 whose survey and interview data met predefined quality-control criteria (see Appendix M). Task 1 (belief-state inference) contains 356 labeled questions, and Task 2 (belief dynamics update) contains 1,386 items, distributed as shown in Table 1.
@@
Belief-State Inf. 108 122 126 356 1,303 1,297 1,302 3,902
Belief Dyn. Update 472 364 550 1,386 3,888 2,610 3,818 10,316
@@ 3.5 Evaluation Protocols
We evaluate belief state inference using accuracy, the proportion of exact matches with ground-truth labels. For belief dynamic update, we report four metrics: (i) accuracy, the proportion of predictions within a tolerance band of the true response ($\pm$1 for 5-point, $\pm$2 for 10-point scales); (ii) mean absolute error (MAE), the average magnitude of deviation from the ground truth, with all responses normalized to a 5-point scale for consistency; (iii) directional accuracy measures whether the predicted belief update matches the ground truth in direction(increase, decrease, or no change); and (iv) unified aggregate score(UAS), following the GLUE (Wang et al., 2018) and SuperGLUE (Wang et al., 2019) scoring paradigm, we derive an overall score by averaging task-specific metrics (accuracy, MAE, directional accuracy) and then taking the unweighted mean across belief-state inference and belief-dynamics update, ensuring balanced comparability across tasks.Formal definitions and computation details are given in Appendix Q.1.
@@
We re-contacted a subset of participants for a short-interval (14-day) test–retest study. Across all sessions, 54 participants contributed data. Of these, 18 completed the retest, and 13 were retained following a demographic consistency check. Belief State Inference yielded an accuracy of 84.49% (SD = 0.36, 95% CI: [80.82, 88.16]). Belief Dynamic Update achieved an accuracy of 85.66% (SD = 1.27, 95% CI: [83.34, 87.70]) and a mean absolute error of 0.67 (SD = 0.83, 95% CI: [0.63, 0.73]). These scores establish a human consistency ceiling, as outlined in Section 3.1, against which model performance can be benchmarked.
@@ 4.1 Baselines
To evaluate the role of agent-like structure in modeling belief reasoning, we further include two agent-style LLM baselines that incorporate memory or retrieval. The first is the Generative Agents baseline, reproduced following the setup of Park et al. (2023), using Qwen2.5-32B-instruct as the base model. The second group includes two variants of retrieval-augmented generation (RAG). The first, RAG, follows the standard setup (Lewis et al., 2020), replacing the original full QA context with the top-$k$ retrieved QA pairs ($k=5$). The second, RAG with Full Context, appends the retrieved QA pairs to the original input, allowing the model to jointly condition on both. This variant evaluates whether retrieval can serve as an auxiliary signal rather than a substitute for agent-specific context.
Detailed settings and full results are provided in Appendix C.
@@ 4.2 Overall Performance
Tables 2 summarize performance. We evaluate using the metrics introduced in Section 3.5. For belief-state inference, best-performing LLMs approach but do not match human accuracy, trailing by 4–6 points. Open-source LLaMA and Qwen rival GPT-4o, while smaller or less aligned models lag significantly. For belief dynamic update, gaps are larger: models frequently mispredict the direction of stance change or fail to adjust reason weights, yielding higher error than human baseline.
@@ 5 Main Findings
Finding 1: More Context Doesn’t Always Help
Model Context Belief-State Inference (Accuracy $\uparrow$) Belief Dynamic Update (Accuracy $\uparrow$) Belief Dynamic Update (MAE $\downarrow$) Health Surv. Zone Avg Health Surv. Zone Avg Health Surv. Zone Avg GPT-4o 5 QAs 68.89±0.51 73.93±1.07 72.22±0.00 71.68 61.48±0.35 67.80±0.45 63.27±0.00 64.19 1.26±0.01 1.19±0.01 1.24±0.00 1.23 10 QAs 67.59±0.17 72.30±0.37 72.06±0.66 70.65 62.29±0.33 64.56±0.39 65.82±0.39 64.22 1.28±0.00 1.26±0.01 1.21±0.01 1.25 20+ QAs 71.11±0.77 75.25±0.90 77.62±0.35 74.66 61.57±0.19 63.74±0.58 64.04±0.20 63.12 1.27±0.01 1.35±0.01 1.26±0.01 1.29 Gemini 2.0 Flash 5 QAs 61.30±0.41 65.25±0.45 65.08±0.00 63.88 59.70±0.09 65.71±0.30 62.25±0.21 62.55 1.36±0.00 1.23±0.00 1.34±0.00 1.31 10 QAs 60.19±0.65 64.75±0.00 71.27±0.35 65.40 58.22±0.28 62.58±0.23 65.20±0.16 61.97 1.37±0.00 1.30±0.01 1.26±0.00 1.31 20+ QAs 63.89±0.00 68.20±0.37 77.78±0.00 69.96 56.23±0.12 62.25±0.37 63.16±0.10 60.55 1.39±0.00 1.33±0.00 1.33±0.00 1.35 Qwen2.5-32B-instr 5 QAs 68.33±1.21 73.11±2.68 65.08±0.56 68.84 59.45±0.55 62.53±0.57 63.38±0.69 61.79 1.36±0.00 1.31±0.01 1.27±0.01 1.31 10 QAs 75.37±4.47 75.90±1.80 67.78±1.06 73.02 58.52±0.74 60.66±0.60 63.02±0.47 60.73 1.41±0.01 1.36±0.01 1.27±0.01 1.35 20+ QAs 77.96±0.41 76.56±0.45 76.98±0.79 77.17 56.82±0.41 57.80±0.25 62.25±0.16 58.96 1.43±0.00 1.43±0.01 1.33±0.01 1.40
Table 3: Question masking / length scaling for both Belief-State Inference and Belief Dynamic Update. Reported as mean $\pm$ std over 5 runs and across domains (healthcare, surveillance, zoning). Best results per column are highlighted in bold.
The core result is that belief-state accuracy rises monotonically with additional dialogue and saturates at 20+ questions, whereas update accuracy peaks at 5–10 questions before declining. This pattern indicates that while longer transcripts provide richer cues for recovering belief states, they also introduce noise that impairs belief updating. The implication is that context length exerts asymmetric effects on the two tasks: recognition benefits from scale, but update is vulnerable to cognitive overload, suggesting that future systems should adaptively select compact spans rather than rely on maximal context.
@@
Finding 2: Personalization Transfer is Harder Than expected
@@
Task Model Health $\rightarrow$ Surv Surv $\rightarrow$ Zone Zone $\rightarrow$ Health Avg Belief-State Inference (Accuracy $\uparrow$) Qwen2.5-32B-instr 53.52±1.52 58.69±1.10 55.08±1.20 55.76 GPT-4o 58.52±0.77 65.41±1.58 51.75±1.03 58.56 Belief Dynamic Update (Accuracy $\uparrow$) Qwen2.5-32B-instr 44.41±0.98 24.84±0.60 30.15±0.35 33.13 GPT-4o 49.49±0.87 42.47±0.50 41.96±0.30 44.64 Belief Dynamic Update (MAE $\downarrow$) Qwen2.5-32B-instr 1.69±0.01 1.94±0.01 2.05±0.01 1.89 GPT-4o 1.55±0.01 1.60±0.00 1.71±0.00 1.62
@@
Table 4: Cross-topic swap test: models are trained with QA context from one domain and evaluated on another domain for the same participant. We report performance for Healthcare $\rightarrow$ Surveillance, Surveillance $\rightarrow$ Zoning, Zoning $\rightarrow$ Healthcare, and their average. Reported as mean $\pm$ std over 5 runs and across domains.
When tested across domains, personalization breaks down: GPT-4o drops to only 58.56% belief-state accuracy on average, far below its within-domain performance, and Qwen2.5-32B fares even worse. The effect is especially pronounced for belief updates, where errors compound across topics. This shows that current models achieve domain-specific fitting rather than generalizable reasoning, highlighting the need to evaluate within-person, cross-domain consistency as a core criterion for robust personalization.
6 Why it Happens: Ablations, Error Analysis and Mitigation Strategies
6.1 Global Prior vs. Full Context
Task Method Accuracy (% $\uparrow$) MAE ($\downarrow$) Healthcare Surveillance Zoning Avg Healthcare Surveillance Zoning Avg Belief-State Inference GPT-4o (No-Context) 55.93±0.01 66.07±0.01 53.49±0.02 58.50±0.02 N/A GPT-4o (Full) 71.11±0.01 75.25±0.01 77.62±0.00 74.66±0.01 N/A Qwen2.5-32B (No-Context) 45.19±0.02 59.02±0.01 48.57±0.01 50.93±0.01 N/A Qwen2.5-32B (Full) 77.96±0.00 76.56±0.00 76.98±0.01 77.17±0.01 N/A Belief Dynamic Update GPT-4o (No-Context) 34.62±0.00 48.21±0.00 36.45±0.00 39.77±0.06 1.77±0.01 1.51±0.01 1.84±0.01 1.71±0.01 GPT-4o (Full) 61.57±0.00 63.74±0.01 64.04±0.00 63.12±0.00 1.27±0.01 1.35±0.01 1.26±0.01 1.29±0.01 Qwen2.5-32B (No-Context) 34.58±0.01 33.90±0.01 27.89±0.00 32.12±0.01 1.89±0.01 1.75±0.01 1.99±0.01 1.88±0.01 Qwen2.5-32B (Full) 56.82±0.00 57.80±0.00 62.25±0.00 58.96±0.00 1.43±0.00 1.43±0.01 1.33±0.01 1.40±0.01
Table 5: Comparison of No-Context (population prior) and Full-Context (with transcripts) settings. Reported as mean $\pm$ std over 5 runs.
Table 5 tests whether models depend only on population priors or can adapt to individual cues. The No-Context setting gives only demographic background, while the Full-Context setting adds transcripts and survey answers. Models show large gains from personalization: for example, GPT-4o’s belief-state inference accuracy rises from 58.50% with only demographic priors to 74.66% with full context, and its belief dynamics update accuracy improves from roughly 38.52% to 63.12%. The consistent gains in Full-Context show that models use personalized information rather than global averages. This highlights that the benchmark can separate real belief inference from reliance on general trends.

### v3

# HugAgent: Benchmarking LLMs for Simulation of Individualized Human Reasoning

@@ Abstract
Simulating human reasoning in open-ended tasks has long been a central aspiration in AI and cognitive science. While large language models now approximate human responses at scale, they remain tuned to population-level consensus, often erasing the individuality of reasoning styles and belief trajectories. To advance the vision of more human-like reasoning in machines, we introduce HugAgent (Human-Grounded Agent Benchmark), which rethinks human reasoning simulation along three dimensions: (i) from averaged to individualized reasoning, (ii) from behavioral mimicry to cognitive alignment, and (iii) from vignette-based to open-ended data. The benchmark evaluates whether a model can predict a specific person’s behavioral responses and the underlying reasoning dynamics in out-of-distribution scenarios, given partial evidence of their prior views. HugAgent adopts a dual-track design: a human track that automates and scales the think-aloud method to collect ecologically valid human reasoning data and a synthetic track for further scalability and systematic stress testing. This architecture enables low-cost, extensible expansion to new tasks and populations. Experiments with state-of-the-art LLMs reveal persistent adaptation gaps, positioning HugAgent as the first extensible benchmark for aligning machine reasoning with the individuality of human thought. The benchmark, along with its complete data collection pipeline and companion chatbot, is open-sourced as HugAgent and TraceYourThinking.
@@ 1 Introduction
Large language models (LLMs) are increasingly used as social simulators—to role-play individuals, build digital twins, and generate synthetic (‘silicon’) samples for testing social and policy ideas(park2023thousand; argyle2023out; xie2024can; jiang2022communitylmprobingpartisanworldviews). These systems promise scalability and accessibility: instead of recruiting thousands of people, researchers and practitioners can use LLMs to approximate human perspectives at scale. Yet because LLMs are pretrained on population-level corpora, they tend to collapse into an “average voice,” capturing consensus patterns while erasing the individuality of personal histories, beliefs, and reasoning styles(wang2025large; santurkar2023whose; durmus2023towards).
@@
Current benchmarks fail to capture this ability, across three key dimensions. 1 Intra-agent vs. inter-agent fidelity. Existing pluralistic alignment benchmarks probe group dynamics and social influence(sorensen2024roadmap), but neglect whether models can faithfully reproduce reasoning within a single agent, which is crucial for identity-consistent modeling. 2 Reasoning traces vs. behavioral outcomes. Large-scale “digital twin” datasets such as Agent Bank (park2023thousand) and Twin-2K-500 (toubia2025twin2k500datasetbuildingdigital) primarily assess static behavioral outcomes, but not the evolving reasoning trajectories of a single individual, which are essential for credible social simulation (li2025positionsimulatingsocietyrequires). 3 Open-ended vs. vignettes. Commonsense and social reasoning benchmarks (e.g., SocialIQA, ATOMIC) often reduce diverse answers to a single ground truth (sap2019atomicatlasmachinecommonsense; sap2019socialiqa). Opinion-oriented datasets likewise emphasize aggregate patterns over individual variation (argyle2023out; santurkar2023whose). Theory-of-Mind style tests typically rely on short vignettes with designer labels (wimmer1983beliefaboutbelifs; chen2024tombenchbenchmarkingtheorymind), which limits ecological validity and overlooks first-person reasoning traces that could serve as a richer gold standard (ying2025benchmark).
@@
Motivated by these gaps, we introduce HugAgent, a benchmark that targets intra-agent fidelity by operationalizing average-to-individual reasoning adaptation as a measurable task. For Dimension 1, HugAgent shifts the granularity from inter-agent to intra-agent fidelity: given a person’s profile and reasoning history, a model must predict both their current belief state and how it would evolve when presented with new counterfactual evidence. In Dimension 2, HugAgent advances beyond static outcomes toward reasoning trajectories. It collects first-person, out-loud self-reports as gold-standard reasoning traces. These traces offer a deeper target for prediction than the choice outcomes or survey responses typically captured in lab experiments. To address Dimension 3, instead of relying on vignette-style benchmarks, HugAgent builds evaluation around open-ended contexts. We curated real-world topics, beginning with socially and politically controversial issues that introduce inherent conflicts. Through sustained follow-up questions, the benchmark probes participants’ deliberate, System 2 style reasoning(kahneman2011thinking; evans2013dual), transforming the dataset from toy settings into complex, open-ended domains. For a broader discussion of prior work on personalization, social reasoning, and user modeling in LLMs, we refer readers to Appendix A.
To build such a benchmark at scale while retaining ecological validity(i.e., the extent to which findings reflect reasoning as it occurs in real-world contexts), HugAgent combines two complementary tracks. A synthetic track provides large, controlled datasets where belief shifts and reasoning paths can be systematically manipulated(yukhymenko2024synthetic; xie2024human). A human-grounded track applies the same protocol to real human participants, yielding data anchored in individual variation(srivastava2023imitationgamequantifyingextrapolating).
@@
  • How well do today’s models perform? We introduce HugAgent, a dual-track benchmark (synthetic + human) that evaluates both Belief State Inference and Belief Dynamics Update. Initial experiments with state-of-the-art LLMs provide baseline results and reveal adaptation gaps. https://anonymous.4open.science/r/HugAgent
@@ 2.1 Formalization
Tasks. (i) Belief State Inference: infer stance/factor polarity from $\mathcal{C}_{i}$. (ii) Belief Dynamics Update: predict stance shifts $\widehat{\Delta\mathbf{s}}_{i}$ given $(\mathcal{C}_{i},\mathcal{I})$. Metrics include accuracy, mean absolute error (MAE), and rank correlation.
@@ 2.2 Theoretical Anchors: Probabilistic and Causal Perspectives
We use normative models as anchors rather than assumptions. (1) Bayesian / PLoT. Idealized revision follows Bayesian conditioning $b_{i}^{\prime}(\mathbf{s})\propto b_{i}(\mathbf{s})\,p(\mathcal{I}\mid\mathbf{s})$, interpreting language as probabilistic evidence over latent stances. We also draw on the Probabilistic Language of Thought (PLoT) framework(goodman2015plot), which extends Bayesian inference to compositional linguistic structures. (2) Structural Causal Models (SCM). Interventions act as $do(\mathcal{I})$ on a causal graph of values/reasons, yielding counterfactual belief shifts $\mathbb{E}[\mathbf{s}\mid do(\mathcal{I})]$. We additionally represent a person’s value–reason structure as a signed directed graph $G_{i}$; we hypothesize that similarity between such graphs (e.g., via graph edit distance or learned embeddings) may predict cross-domain transfer, motivating our Hypothesis H2 and left as future work. Human reasoning deviates from these ideals; the anchors provide principled baselines for analysis.
@@ 3 HugAgent Benchmark
Figure 2: Two benchmark tasks. Task 1 (Belief State Inference) infers stance and reasons from prior context; Task 2 (Belief Dynamics Update) predicts stance shifts and its reasoning under new evidence.
@@ 3.1 Design Principles
  • Open-ended but deeper. Emphasize depth over breadth: semi-structured dialogue with targeted follow-ups surfaces individuality while avoiding over-scaffolding; ground truth comes from self-reports (kvale1996interviews; srivastava2023imitationgamequantifyingextrapolating; park2023thousand).
  • Two observable proxies. We evaluate (i) belief state inference and (ii) belief dynamics update—tractable targets that avoid requiring exact trace imitation; cf. proxy-label benchmarks (geva2021didaristotleuselaptop; ho2022wikiwhyansweringexplainingcauseandeffect; guerdan2023groundlesstruthcausalframework).
@@
  • Human ceiling. Test–retest reliability defines the upper bound, aligning with psychology standards (nunnally1994psychometric; cronbach1970essentials) and recent large-scale simulations (toubia2025twin2k500datasetbuildingdigital; park2023thousand).
@@ 3.2 Task Definition
We instantiate two tasks: (i) Belief State Inference: predict $(s_{t},\mathbf{w}_{t})$ from prior responses. (ii) Belief Dynamics Update: predict $(s_{t+1},\mathbf{w}_{t+1})$ given $(b_{t},e)$. Figure 2 shows concrete examples of both tasks in HugAgent.
Figure 3: HugAgent benchmark pipeline. Inputs (demographics, questionnaires, and transcripts) flow through two components: a questionnaire that provides demographic anchors, stance baselines, and counterfactual updates, and a semi-structured chatbot that elicits individualized reasoning. Together these elements define two benchmark tasks: (i) belief state inference—recovering stance and factor polarity from context, and (ii) belief dynamics update—predicting stance shifts and reweighting under new evidence. (A) The chatbot maintains a causal belief network of factors, used to identify the most critical nodes and edges for follow-up. (B) A question generator derives targeted, context-specific probes from this network to structure the dialogue.
@@ 3.3 Building HugAgent: Scalable Elicitation of Individual Reasoning
HugAgent is built through a two–stage pipeline (Figure 3).
1. The questionnaire stage collects demographics, a baseline stance ($s_{t}$, 1–10), reason weights ($\mathbf{w}_{t}$, 1–5), and counterfactual interventions with updated stances and reasons. These structured responses provide gold labels for Belief Dynamics Update and serve as anchors for aligning free text to factors.
2. The chatbot stage elicits 8–20 question–answer pairs through semi-strutctured interview, combining open-ended elaborations (Context QAs) with concise polarity judgments (GT QAs). This setup captures both participants’ belief, reasoning styles and explicit preferences on each decision factor. Each transcript thus supports both benchmark tasks: Belief State Inference (using Context and GT QAs) and Belief Dynamics Update (using Context QAs, questionnaire responses upon interventions). Survey-provided updates are never revealed in dialogue, preventing leakage.
@@ 3.4 Dataset Statistics
From over 120 participants, we retained 54 whose survey and interview data met predefined quality-control criteria (see Appendix M). Task 1 (belief state inference) contains 356 labeled questions, and Task 2 (belief dynamics update) contains 1,386 items, distributed as shown in Table 1.
@@
Belief State Inference 108 122 126 356 1,303 1,297 1,302 3,902
Belief Dynamics Update 472 364 550 1,386 3,888 2,610 3,818 10,316
@@ 3.5 Evaluation Protocols
We evaluate belief state inference using accuracy, the proportion of exact matches with ground-truth labels. For belief dynamics update, we report four metrics: (i) accuracy, the proportion of predictions within a tolerance band of the true response ($\pm$1 for 5-point, $\pm$2 for 10-point scales); (ii) mean absolute error (MAE), the average magnitude of deviation from the ground truth, with all responses normalized to a 5-point scale for consistency; (iii) directional accuracy measures whether the predicted belief update matches the ground truth in direction(increase, decrease, or no change); and (iv) average to individual (ATI) score, following the GLUE (wang2018glue) and SuperGLUE (wang2019superglue) scoring paradigm, we derive an overall score via hierarchical aggregation, where normalized task-specific metrics (accuracy, MAE, directional accuracy) are combined into belief dynamics update and averaged with belief state inference, with human and random guess performance used as upper and lower bounds for normalization. Formal definitions and computation details are given in Appendix Q.1.
@@
We re-contacted a subset of participants for a short-interval (14-day) test–retest study. Across all sessions, 54 participants contributed data. Of these, 18 completed the retest, and 13 were retained following a demographic consistency check. Belief State Inference yielded an accuracy of 84.84% (SD = 8.90, 95% CI: [80.00, 89.68]). Belief Dynamics Update achieved an accuracy of 85.66% (SD = 7.66, 95% CI: [80.91, 90.40]) and a mean absolute error of 0.68 (SD = 0.20, 95% CI: [0.55, 0.80]). The directional accuracy was 88.92% (SD = 9.89, 95% CI: [82.09, 96.24]). These scores establish a human consistency ceiling, as outlined in Section 3.1, against which model performance can be benchmarked.
@@ 4.1 Baselines
To evaluate the role of agent-like structure in modeling belief reasoning, we further include two agent-style LLM baselines that incorporate memory or retrieval. The first is the Generative Agents baseline, reproduced following the setup of park2023thousandpeople, using Qwen2.5-32B-instruct as the base model. The second group includes two variants of retrieval-augmented generation (RAG). The first, RAG, follows the standard setup (lewis2020retrieval), replacing the original full QA context with the top-$k$ retrieved QA pairs ($k=5$). The second, RAG with Full Context, appends the retrieved QA pairs to the original input, allowing the model to jointly condition on both. This variant evaluates whether retrieval can serve as an auxiliary signal rather than a substitute for agent-specific context. Detailed settings and full results are provided in Appendix C.
@@ 4.2 Overall Performance
Tables 2 summarize performance. We evaluate using the metrics introduced in Section 3.5. For belief state inference, best-performing LLMs approach but do not match human accuracy, trailing by 4–6 points. Open-source LLaMA and Qwen rival GPT-4o, while smaller or less aligned models lag significantly. For belief dynamics update, gaps are larger: models frequently mispredict the direction of stance change or fail to adjust reason weights, yielding higher error than human baseline.
@@ 5 Main Findings
Finding 1: Preserving Identity Across Domains is Harder Than Expected
@@
Task Model Health $\rightarrow$ Surv Surv $\rightarrow$ Zone Zone $\rightarrow$ Health Avg Belief State Inference (Accuracy % $\uparrow$) Qwen2.5-32B-instr. 53.52±1.52 58.69±1.10 55.08±1.20 55.76±2.54 GPT-4o 58.52±0.77 65.41±1.58 51.75±1.03 58.56±5.88 Belief Dynamics Update (Accuracy % $\uparrow$) Qwen2.5-32B-instr. 44.41±0.98 24.84±0.60 30.15±0.35 33.13±8.58 GPT-4o 49.49±0.87 42.47±0.50 41.96±0.30 44.64±3.60 Belief Dynamics Update (MAE $\downarrow$) Qwen2.5-32B-instr. 1.69±0.01 1.94±0.01 2.05±0.01 1.89±0.16 GPT-4o 1.55±0.01 1.60±0.00 1.71±0.00 1.62±0.07
@@
Table 3: Cross-domain swap test: models are trained with QA context from one domain and evaluated on another domain for the same participant. We report performance for Healthcare $\rightarrow$ Surveillance, Surveillance $\rightarrow$ Zoning, Zoning $\rightarrow$ Healthcare, and their average. Reported as mean $\pm$ std over 5 runs.
To evaluate whether models can generalize a person’s contextual information across domains, we conduct a cross-domain swap test. Each model is given QA context from one domain and evaluated on another domain for the same participant (for example, using context from Healthcare and testing on Surveillance). Based on Table 2, we selected GPT-4o and Qwen2.5-32B-instruct as representative models from two categories: closed-source SOTA and strong open-source baseline.
Under cross-domain transfer, model performance degrades substantially compared to within-domain evaluation. GPT-4o achieves an average of 58.56% on belief state inference, compared to its within-domain score of 74.66%, while Qwen2.5-32B-instruct drops from 77.17% to 55.76%. In the Belief Dynamics Update task, GPT-4o declines from 63.11% to 44.64%. The degradation is most pronounced in belief dynamics update, where compounding errors accumulate across domains. These findings suggest that current models depend heavily on domain-specific linguistic and contextual cues, resulting in limited cross-domain reasoning transfer. This underscores the importance of evaluating within-person, cross-domain consistency as a key indicator of robust generalization, and suggests that improving transferability requires focusing on essential, domain-relevant context rather than surface-level correlations.
Finding 2: More Context Doesn’t Always Help